OpenStudy (anonymous):

Evaluate the function f (x) = 4 • 7x for x= -1 and x = 2. i know how to write out the equation for it but 7^x is throwing me off. do i change it to '2' or do i solve it, i really need help, i will give medals

OpenStudy (bibby):

\(\large a^{-x}=\frac{1}{a^x}\)

OpenStudy (anonymous):

i got f(-1)=4(-1)+7^x

OpenStudy (bibby):

what is 7^-1?

OpenStudy (anonymous):

-49?

OpenStudy (bibby):

follow the property I wrote in the first post. \(\large7^{-4}=\frac{1}{7^4}\\\large5^{-7}=\frac{1}{5^7}\)